OpenStudy (anonymous):

A rectangular garden has a perimeter of 130 feet. First find an equation for the area of the rectangle as a function of the width, then determine the length and width of the rectangle which provide the maximum area. Area = A(w) =

OpenStudy (bahrom7893):

P = 2(w+l) = 130, so w+l = 65

OpenStudy (bahrom7893):

Area = w*l, but w+l = 65, or l = 65-w. Plug that into the area's equation: A = w*(65-w)

OpenStudy (bahrom7893):

Maximize area: A = -w^2 + 65w A' = -2w+65 = 0 -2w=-65 w = 32.5

OpenStudy (bahrom7893):

l = 65-32.5 = 32.5

OpenStudy (bahrom7893):

Max area = 32.5^2